OpenStudy (deadshot):

Given Triangle ABC with a = 24, Angle A=101 degrees , and Angle B=36 degrees, find b. Round the sine value to the nearest thousandth and the answer to the nearest hundredth.

OpenStudy (anonymous):

can you draw a picture???

OpenStudy (mathstudent55):

Do you know the law of sines?

OpenStudy (mathstudent55):

When you have one side and the opposite angle given, you can use the law of sines. \( \dfrac{a}{\sin A} = \dfrac{b}{\sin B} = \dfrac{c}{\sin C} \) In your case, use the first two fractions and use a = 36, A = 101 deg, B = 36 deg. Solve for b.

OpenStudy (deadshot):

I don't know the law of sines, can you help me?

OpenStudy (mathstudent55):

I just did.